How do I work these kind of problems? 2/8=V/2
Answer by TimothyLamb(4379)

2/8 = V/2
(2/8)2 = (V/2)2
4/8 = (2V/2)
1/2 = (2/2)V
1/2 = V
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php